OpenStudy (anonymous):

convert 0.004 g/(min.)(m^3)cm to lbm/(hr)(ft^3)

OpenStudy (anonymous):

can you help me @ganeshie8 and @Michele_Laino :) i can do it if the given is g/(min.)(m^3) ,but there's a cm in the denominator :/

OpenStudy (michele_laino):

is your measure like this? \[\Large 0.04\frac{g}{{\min \times {m^3} \times cm}}\]

OpenStudy (anonymous):

yes :)

OpenStudy (michele_laino):

"lbm" stands for?

OpenStudy (anonymous):

pound (mass)

OpenStudy (michele_laino):

g stands for grams? right?

OpenStudy (anonymous):

yes :)

OpenStudy (michele_laino):

please wait, I'm working on your question

OpenStudy (anonymous):

it's okay and take your time, thank you for helping me :)

OpenStudy (michele_laino):

please note that, in the first unit measure, we divide by a length^4 whereas in the second unit of measure we divide by a length^3

OpenStudy (michele_laino):

since the second unit of measure is: \[\Large \frac{{{\text{lbm}}}}{{{\text{hour}} \cdot {\text{ fee}}{{\text{t}}^{\text{3}}}}}\]

OpenStudy (anonymous):

uhmm i'm sorry I didn't quite get it about dividiing by a length^4 part, can you show it to me. i'm really sorry :(

OpenStudy (michele_laino):

the first unit of measure is: \[\Large \frac{{{\text{grams}}}}{{{\text{min }} \times {{\text{m}}^{\text{3}}}{\text{ }} \times {\text{cm}}}}{\text{ = }}\frac{{{\text{mass}}}}{{{\text{time }} \times {\text{lengt}}{{\text{h}}^{\text{4}}}}}\]

OpenStudy (michele_laino):

m stands for meter, right?

OpenStudy (anonymous):

yes :)

OpenStudy (michele_laino):

so: m^3*cm = length^4

OpenStudy (anonymous):

\[4x10^-5 kg/(\min)(m^4) \times 60 \min./1hr \times 1lbm/0.4536 kg \times (0.3048m)^3/1ft.^3\]

OpenStudy (michele_laino):

so, there is not cm at the denominator?

OpenStudy (anonymous):

oops sorry i'll do it again i forgot to convert cm to m. am I doing it right?

OpenStudy (anonymous):

nope, it says on our worksheet that it's placed in the denominator :(

OpenStudy (michele_laino):

if the dimensions are not equal, we can not change from one unit of measure to another

OpenStudy (michele_laino):

maybe there is m^2 instead of m^3?

OpenStudy (anonymous):

i'll just ask my professor about this quetion, maybe it's a typographical error, it would be easier if cm is not present in the question :)))

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

thank you so much for helping me :D

OpenStudy (michele_laino):

:)